Ask
your own question, for FREE!
Mathematics
7 Online
You roll 2 dice. What is the probability that the sum of the dice is odd or 1 die shows a 5? A 6 X 6 table of dice outcomes will help you to answer this question.
Still Need Help?
Join the QuestionCove community and study together with friends!
The sum of numbers can be either even or odd.... The probability of having a sum as odd is 1/2. Probability of none of the dices showing a four is (5/6)*(5/6) = 25/36 Thus probability of at least one four is 1 - 25/36 = 11/36 Total probability = 1/2 + 11/36 = 29/36
1/6 would be the correct answer
nvm, it would be 23/36
Can't find your answer?
Make a FREE account and ask your own questions, OR help others and earn volunteer hours!
Join our real-time social learning platform and learn together with your friends!
Join our real-time social learning platform and learn together with your friends!
Latest Questions
VERNEISH:
Two objects are moving along separate linear paths where each path is described by position, d, and time, t.
Maci:
what are the best stock apps for people who want to start investing?
Twaylor:
genuienly how do i get better self control i keep watcing youtube and theres this
IzzyCookie:
|dw:1792074056138:dw| Uhm, I don't know where to ask this, but I'm in college welding at my community college, and I need help learning vertical welding for
thehearken:
not sure if this is considered spam, im sorry if it is, but does anyone have tips
17 hours ago
1 Reply
0 Medals
5 days ago
3 Replies
2 Medals
1 week ago
3 Replies
0 Medals
3 days ago
9 Replies
1 Medal
1 week ago
2 Replies
1 Medal